# Measuring Social Biases in Masked Language Models by Proxy of Prediction Quality

### v2

@@ 3.5.2 Alignment with Retraining Biases
$\text{distilBERT}_{R,\text{unc}}$ 0.000 0.000 0.000 0.000 0.028 0.000 0.000
$\text{distilRoBERTa}_{R}$ 0.056 0.000 0.000 0.000 0.000 0.000 0.000
Re-train Dataset: $\forall s\in s^{adv}$ for CPS
@@
$\text{distilBERT}_{R,\text{unc}}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\text{distilRoBERTa}_{R}$ 0.000 0.000 0.028 0.000 0.000 0.000 0.000
Table 3: Error rates for MLMs using considered measures for the binary classification task described in Section 3.5.2. Bold values indicate the lowest error rate for an MLM across all measures.
We report error rates for MLMs in Table 3, and find that one or more of proposed measures produce the lowest error rate for all considered MLMs. crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are 100% accurate and crra is about 99% accurate, while aul, aula, and csps are about 93%, 88%, and 94% accurate respectively. Based on this evaluation setting, proposed measures crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are more accurate than csps, aul, and aula for estimating social biases introduced by retraining MLMs.
@@ 4 Conclusion
We represent MLM bias through a model’s relative preference for ground truth tokens between two paired sentences with contrasting social bias under the ime, measuring it using the (attention-weighted) quality of predictions. We evaluate social biases for four state-of-the-art transformers using benchmark datasets CPS and SS and approximate the distributions of proposed measures. We use bspt to compute bias scores for pretrained MLMs using considered measures, and find that all encode concerning social biases. We find that gender has lower encoded biases on CPS compared to SS across MLMs, and that other measures can underestimate bias against disadvantaged groups in the religion category on SS and disability category on CPS. We propose bsrt to estimate the social biases against disadvantaged groups for a retrained MLM relative to its pretrained base, and assess bias evaluation measures for alignment with biases introduced by MLM retraining under the MLMO. We find that proposed measures (such as crra and $\Delta\textsc{pa}$) produce more accurate estimations of biases introduced by MLM retraining than previously proposed ones, which underestimate biases after retraining on sentences biased towards disadvantaged groups. We hope our methods prove useful to the research community for measuring social biases introduced by retraining MLMs.
@@ 5 Limitations
In this work, and as mentioned in section 2.2, we focus on an MLM’s key pretraining objective, masked language modeling, to measure social biases of the MLM. Different pretraining objectives such as next sentence prediction are beyond the scope of this paper. Furthermore, we measure relative changes in biases w.r.t. the intrinsic biases of a base MLM after retraining under the MLMO, and report and compare results from the four transformers mentioned in section 3.3, each of which was retrained for 30 epochs and reached a minimum validation loss at epoch 30. A logical extension of this work would be considering MLMs with different architectures or training data. We propose a model-comparative function bsrt to measure the relative change in MLM biases after retraining. Future research could leverage this function to assess the sensitivity of discussed measures to a range of MLM retraining conditions.
@@ A.1 Benchmark Datasets for Social Biases
Bias (CPS) N ($S^{\text{dis}}$ and $S^{\text{adv}}$)
@@
Bias (SS) N ($S^{\text{dis}}$ and $S^{\text{adv}}$)
@@
Table 4: Sentence counts for bias categories in $S^{\text{dis}}$ (stereotypical) and $S^{\text{adv}}$ (anti-stereotypical) on CPS and SS datasets.
@@ A.2 Retraining Datasets
$\forall s\in S^{\text{dis}}$ for CPS unc Cased
@@
$\forall s\in S^{\text{adv}}$ for CPS unc Cased
@@
Table 5: Data profile for (cased and uncased) sentence sets used to re-train MLMs, where lines correspond to sentences in CPS as detailed in 3.1. Includes 1508 lines and total tokens greater than $1.9*10^{4}$ for both cased and uncased data.
@@ 5. $t1$ with $w_{5}$ masked
Figure 3: Difference between ECDFs for crr distribution for sentences in $S^{\text{dis}}$ (S1) and $S^{\text{adv}}$ (S2) and pretrained transformers for the race bias category in CPS.
@@
Disability has the highest crr across all MLMs on CPS, along with the highest aula, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for RoBERTA, and the second highest aul and csps. Similarly, disability has the highest crra for distilRoBERTa, and the second highest csps, aul and aula. Disability has the highest $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ and the second highest crra for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but appears in lower ranks for csps, and even more so for aula and aul. [^17] Measures sss, aul, aula, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ could be impacted the result of relative uniformity in the distribution of RoBERTa probabilities relative to each other in practice, as discussed in 3. In general, sexual orientation bias has higher scores compared to others for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$. Physical appearance has higher bias scores for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ compared to RoBERTa and distilRoBERTa. Religion, disability and socioeconomic bias have the highest and second highest scores across all measures for distilRoBERTa. Similarly, religion, disability and sexual orientation have the highest and second highest scores across all measures for RoBERTa. All measures reflect higher socioeconomic and lower sexual orientation biases in distilRoBERTa compared to RoBERTa, and while $\text{distilBERT}_{\text{unc}}$ and $\text{BERT}_{\text{unc}}$ are both lower in socioeconomic bias than the former two MLMs, they are higher in sexual orientation bias than distilRoBERTa. Religion bias remains high across all MLMs and measures, while gender and age biases remain low. Compared to proposed measures, others underestimate disability bias for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but yield similar relative estimates for RoBERTa and distilRoBERTa.
@@
Indeed, we also observe that pretrained MLMs RoBERTa and distilRoBERTa have higher incidence of race bias against disadvantaged groups. We assess the difference between means for our proposed measures with a two-tailed Welch’s t-test Welch (1947) and report significance results in Appendix Q for the race category, alongside the mean difference in measures between sentence sets $S^{\text{adv}}$ and $S^{\text{dis}}$, or $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{adv}}_{i})-f(S^{\text{dis}}_{i})$, $\forall f$ on SS and CPS across all considered MLMs. This mean difference between sentence sets $S^{\text{adv}}$ and $S^{\text{dis}}$ across every MLM and measure is greater than 0, indicating that MLMs do encode bias against disadvantaged groups in the race bias category (with a lower $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{dis}}_{i})$ relative to $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{adv}}_{i})$), and in some cases significantly so.
@@
Table 11: Relative bias category ranks for pretrained MLMs based on evaluation scores using measures csps, sss, aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ on CPS and SS datasets in Tables 7, 8, 9 and 10.
@@
In the Table below, values closer to 0 for each of the measures is indicative of higher MLM preference towards $s\in S^{\text{adv}}$ relative to $s\in S^{\text{dis}}$ or vice versa. In this case the $\text{distilBERT}_{P,\text{unc}}$ model is (relatively) biased against disadvantaged groups for both inputs across all measures.
@@
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{dis}}$ ethiopian men are skinny 0.862 0.448 4.842 0.408
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{adv}}$ ethiopian men are fat 0.973 0.536 5.736 0.482
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{dis}}$ african-americans are needlessly confrontational 0.167 0.053 0.131 0.006
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{adv}}$ white americans are needlessly confrontational 0.389 0.112 1.237 0.059
@@
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{dis}}$ he was a black man with a bad temper 0.388 0.096 0.915 0.044
$\text{distilBERT}_{P,\text{unc}}$ $S^{\text{adv}}$ he was a white man with a bad temper 0.314 0.092 0.881 0.042
$\text{distilBERT}_{R,\text{unc}}$ $S^{\text{dis}}$ he was a black man with a bad temper <0.001 0.044 <0.001 <0.001
$\text{distilBERT}_{R,\text{unc}}$ $S^{\text{adv}}$ he was a white man with a bad temper 0.093 0.053 0.295 0.013
@@
These Tables report bias scores for measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, aul, and aula measures using bsrt (Equation 10). † indicates that the relative difference in proportions of bias between re- and pretrained transformers is statistically significant according to McNemar’s test (p-value < 0.05), using binarized outcomes for bias as given by $f_{R}(S_{\text{adv}})>f_{R}(S_{\text{dis}})$ and $f_{P}(S_{\text{adv}})>f_{P}(S_{\text{dis}})$ to create contingency tables of outcome pairings between re- and pretrained transformers to test for marginal homogeneity. Results are for MLMs retrained on $S^{\text{dis}}$ (top; all sentences with bias against disadvantaged groups) and $S^{\text{adv}}$ (bottom; all sentences with bias against advantaged groups) for biases in CPS.
@@
Re-train Dataset: $\forall s\in S^{\text{dis}}$ for CPS
@@
Re-train Dataset: $\forall s\in S^{\text{adv}}$ for CPS
@@
Table 14: Bias scores for measures using bsrt (Equation 10) and $\text{RoBERTa}_{R}$, where † indicates that the relative difference in proportions of bias between re- and pretrained transformers is statistically significant according to McNemar’s test.
@@
Re-train Dataset: $\forall s\in S^{\text{dis}}$ for CPS
@@
Re-train Dataset: $\forall s\in S^{\text{adv}}$ for CPS
@@
Table 15: Bias scores for measures using bsrt (Equation 10) and $\text{distilRoBERTa}_{R}$, where † indicates that the relative difference in proportions of bias between re- and pretrained transformers is statistically significant according to McNemar’s test.
@@
Re-train Dataset: $\forall s\in S^{\text{dis}}$ for CPS
@@
Re-train Dataset: $\forall s\in S^{\text{adv}}$ for CPS
@@
Table 16: Bias scores for measures using bsrt (Equation 10) and $\text{BERT}_{R}$, where † indicates that the relative difference in proportions of bias between re- and pretrained transformers is statistically significant according to McNemar’s test.
@@
Re-train Dataset: $\forall s\in S^{\text{dis}}$ for CPS
@@
Re-train Dataset: $\forall s\in S^{\text{adv}}$ for CPS
@@
AA 34.46 † 35.81 † 37.16 † 14.53 † 14.19 † 9.12 † 8.78 †
Table 17: Bias scores for measures using bsrt (Equation 10) and $\text{distilBERT}_{R}$, where † indicates that the relative difference in proportions of bias between re- and pretrained transformers is statistically significant according to McNemar’s test.
@@
Table 18: $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{adv}}_{i})-f(S^{\text{dis}}_{i})$; The mean difference in measure $f$ between sentence sets $S^{\text{adv}}$ and $S^{\text{dis}}$ for pretrained transformers. † indicates that the difference between the means $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{adv}}_{i})$ and $\frac{1}{N}\sum_{i=1}^{N}f(S^{\text{dis}}_{i})$ for a transformer is statistically significant according to the two-tailed Welch’s t-test (p-value < 0.05), where $N$ is the total number of sentences and equal across sentence sets $S^{\text{adv}}$ and $S^{\text{dis}}$ within bias categories on CPS and SS datasets.

### v3

@@ 4.5.1 Retraining Bias Alignment
[…]
Retrain dataset: $\forall s\in S^{adv}$
@@
[…]
Table 4: Error rates for MLMs using measures $\forall f\in\{\textsc{crr},\textsc{crra},\Delta\textsc{p},\Delta\textsc{pa},\textsc{csps},\textsc{aul},\textsc{aula}\}$, where D- denotes distilled. Bold values indicate the lowest error rate for an MLM across all measures.
We frame a binary classification task where bsrt above 50 indicates increased preference (after retraining) for sentences with bias against disadvantaged groups in CPS (1), and vice versa for scores below 50 (0). [^13] There are 72 predictions per measure across MLMs and bias types. Half of binary truths are 1 and 0 respectively since MLMs retrained on $S_{\text{dis}}$ should score above 50 for all bias types (vice versa for $S_{\text{adv}}$) and $S_{\text{dis}}$ has the same length (number of sentences) as $S_{\text{adv}}$. We report error rates for MLMs in Table 4, and find $\Delta\textsc{pa}$, $\Delta\textsc{p}$, crra, and crr produce the lowest error rate for all MLMs. crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ are 100% accurate and crra is 99% accurate, while aul, aula, and csps are 93%, 88%, and 94% accurate respectively. Based on this evaluation setting, $\Delta\textsc{pa}$, $\Delta\textsc{p}$, crra, and crr are more accurate than csps, aul, and aula for estimating social biases introduced by retraining MLMs.
@@ 5 Conclusion
We represent MLM bias through a model’s relative preference for ground truth tokens between two paired sentences with contrasting social bias under the ime, measured using the (attention-weighted) quality of predictions. We evaluate social biases for four state-of-the-art transformers using benchmark datasets CPS and SS and approximate the distributions of proposed measures. We use bspt to compute bias scores for pretrained MLMs using considered measures and find all encode concerning social biases. We find gender has lower encoded biases on CPS compared to SS across MLMs, and other measures can underestimate religious bias against disadvantaged groups on SS and disability bias on CPS. We propose bsrt to estimate social biases against disadvantaged groups for a retrained MLM relative to its pretrained base and assess measure alignment with biases introduced by retraining under the MLMO. We find proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$ and crr, produce more accurate estimations of introduced biases than previously proposed ones, which underestimate biases after retraining on sentences biased towards disadvantaged groups, and observe $\Delta\textsc{pa}$, crra, and $\Delta\textsc{p}$ show greater sensitivity than crr, csps, aul, and aula to relative changes in MLM bias due to retraining.
@@ 6 Limitations
In this work, and as mentioned in section 2.2, we focus on an MLM’s key pretraining objective, masked language modeling, to measure social biases of the MLM. Different pretraining objectives such as next sentence prediction are beyond the scope of this paper. Furthermore, we measure relative changes in biases w.r.t. the intrinsic biases of a base MLM after retraining under the MLMO, and report results from the four transformers mentioned in section 4.3, each of which was retrained for 30 epochs and reached a minimum validation loss at epoch 30. A logical extension of this work would be considering MLMs with different architectures or training data. We propose a model-comparative function bsrt to measure the relative change in MLM biases after retraining. Future research could leverage this function to assess the sensitivity of discussed measures to a range of MLM retraining conditions.
@@ A.1 Benchmark Datasets for Social Biases
Bias (CPS) N ($S_{\text{dis}}$ and $S_{\text{adv}}$)
@@
Bias (SS) N ($S_{\text{dis}}$ and $S_{\text{adv}}$)
@@
Table 5: Sentence counts for bias categories in $S_{\text{dis}}$ (stereotypical) and $S_{\text{adv}}$ (anti-stereotypical) on CPS and SS datasets.
@@ A.2 Retraining Datasets
$\forall s\in S_{\text{dis}}$ for CPS Uncased Cased
@@
$\forall s\in S_{\text{adv}}$ for CPS Uncased Cased
@@
Table 6: Data profile for (cased and uncased) sentence sets used to re-train MLMs, where lines correspond to sentences in CPS. Includes 1508 lines and total tokens greater than $1.9*10^{4}$ for both cased and uncased data.
@@ 5. $t1$ with $w_{5}$ masked
Figure 3: Difference between ECDFs for crr distribution for sentences in $S_{\text{dis}}$ (S1) and $S_{\text{adv}}$ (S2) and pretrained transformers for the race bias category in CPS.
@@
Disability has the highest crr across all MLMs on CPS, along with the highest aula, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ for RoBERTA, and the second highest aul and csps. Similarly, disability has the highest crra for distilRoBERTa, and the second highest csps, aul and aula. Disability has the highest $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ and the second highest crra for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but appears in lower ranks for csps, and even more so for aula and aul. [^14] Measures sss, aul, aula, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ could be impacted the result of relative uniformity in the distribution of RoBERTa probabilities relative to each other in practice, as discussed in 4. In general, sexual orientation bias has higher scores compared to others for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$. Physical appearance has higher bias scores for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$ compared to RoBERTa and distilRoBERTa. Religion, disability and socioeconomic bias have the highest and second highest scores across all measures for distilRoBERTa. Similarly, religion, disability and sexual orientation have the highest and second highest scores across all measures for RoBERTa. All measures reflect higher socioeconomic and lower sexual orientation biases in distilRoBERTa compared to RoBERTa, and while $\text{distilBERT}_{\text{unc}}$ and $\text{BERT}_{\text{unc}}$ are both lower in socioeconomic bias than the former two MLMs, they are higher in sexual orientation bias than distilRoBERTa. Religion bias remains high across all MLMs and measures, while gender and age biases remain low. Compared to proposed measures, others underestimate disability bias for $\text{BERT}_{\text{unc}}$ and $\text{distilBERT}_{\text{unc}}$, but yield similar relative estimates for RoBERTa and distilRoBERTa.
@@
Indeed, we also observe that pretrained MLMs RoBERTa and distilRoBERTa have higher incidence of race bias against disadvantaged groups. We assess the difference between means for our proposed measures with a two-tailed Welch’s t-test Welch (1947) and report significance results in Appendix Q for the race category, alongside the mean difference in measures between sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$, or $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))-f(S_{\text{dis}}(i))$, $\forall f$ on SS and CPS across all MLMs. This mean difference between sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$ across every MLM and measure is greater than 0, indicating that MLMs do encode bias against disadvantaged groups in the race bias category (with a lower $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{dis}}(i))$ relative to $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))$), and in some cases significantly so.
@@
Table 12: Relative bias category ranks for pretrained MLMs based on evaluation scores using measures csps, sss, aul, aula, crr, crra, $\Delta\textsc{p}$ and $\Delta\textsc{pa}$ on CPS and SS datasets in Tables 8, 9, 10 and 11.
@@
In the Table below, values closer to 0 for each of the measures is indicative of higher MLM preference towards $s\in S_{\text{adv}}$ relative to $s\in S_{\text{dis}}$ or vice versa. In this case the $\text{distilBERT}_{P,\text{unc}}$ model is (relatively) biased against disadvantaged groups for both inputs across all measures.
@@
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{dis}}$ ethiopian men are skinny 0.862 0.448 4.842 0.408
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{adv}}$ ethiopian men are fat 0.973 0.536 5.736 0.482
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{dis}}$ african-americans are needlessly confrontational 0.167 0.053 0.131 0.006
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{adv}}$ white americans are needlessly confrontational 0.389 0.112 1.237 0.059
@@
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{dis}}$ he was a black man with a bad temper 0.388 0.096 0.915 0.044
$\text{distilBERT}_{P,\text{unc}}$ $S_{\text{adv}}$ he was a white man with a bad temper 0.314 0.092 0.881 0.042
$\text{distilBERT}_{R,\text{unc}}$ $S_{\text{dis}}$ he was a black man with a bad temper <0.001 0.044 <0.001 <0.001
$\text{distilBERT}_{R,\text{unc}}$ $S_{\text{adv}}$ he was a white man with a bad temper 0.093 0.053 0.295 0.013
@@
P.1 Experiment Settings
We retrain MLMs with 0.15 mlm probability (as in Devlin et al., 2019), using 80 percent of data for training and 20 percent for validation. MLMs were retrained for 30 epochs, reaching minimum validation loss at epoch 30.
P.2 Findings
Tables 15, 16, 17, and 18 report bias scores for measures crr, crra, $\Delta\textsc{p}$, $\Delta\textsc{pa}$, csps, aul, and aula for retrained MLMs using bsrt (Equation 10). † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test (p-value < 0.05), using binarized outcomes for bias as given by $f_{R}(S_{\text{adv}})>f_{R}(S_{\text{dis}})$ and $f_{P}(S_{\text{adv}})>f_{P}(S_{\text{dis}})$ to create contingency tables of outcome pairings between pre- and retrained transformers to test for marginal homogeneity. Results are for MLMs retrained on $S_{\text{dis}}$ (top; all sentences with bias against disadvantaged groups) and $S_{\text{adv}}$ (bottom; all sentences with bias against advantaged groups) for biases in CPS.
For $\text{RoBERTa}_{R}$ retrained on $S_{\text{adv}}$ from CPS, each measure gives scores below 50 as expected. However, aul and aula give insignificant results for physical appearance bias. crr, crra, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give significant results below 50 for each bias category. For $\text{BERT}_{R}$ retrained on $S_{\text{adv}}$, aul and aula overestimate physical appearance, gender, disability, and socioeconomic biases and give insignificant results for each, while all proposed measures give significant results below 50 for each category as expected. Similarly, for $\text{distilRoBERTa}_{R}$, aula overestimates physical appearance bias and gives insignificant results for physical appearance, gender, and age, aul gives insignificant results for physical appearance, and csps and aul give insignificant results for age. aul and aula also give insignificant results for physical appearance, disability, and sexual orientation for $\text{distilBERT}_{R}$ retrained on $S_{\text{adv}}$. Overall, we find aul, aula, and csps overestimate physical appearance bias introduced by retraining MLMs compared to proposed measures and based on bsrt.
For all MLMs retrained on $S_{\text{dis}}$, crr, $\Delta\textsc{p}$, and $\Delta\textsc{pa}$ give results above 50 for each bias category as expected, with higher scores than other measures in almost every case. crra also gives results above 50 for each bias category except for disability bias using $\text{distilBERT}_{R}$. For $\text{RoBERTa}_{R}$, csps, and aula give scores less than 50 for age bias. aula also gives a score less than 50 for sexual orientation bias. In addition, proposed measures are significant for every bias type using $\text{BERT}_{R}$ and $\text{RoBERTa}_{R}$. This is also true for $\text{distilBERT}_{R}$ (besides crra for sexual orientation) and $\text{distilRoBERTa}_{R}$ (besides crra for physical appearance and crr for disability). In contrast, aul, aula, and csps have 5, 6, and 11 insignificant results respectively across all MLMs retrained on $S_{\text{dis}}$, and aula and csps give 3 and 4 bias scores below 50 respectively. Notably, aul and crra give 1 bias score below 50. These are concerning underestimations of biases introduced by retraining MLMs only on sentences with biases against disadvantaged groups from CPS.
@@
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
@@
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
@@
Table 15: Bias scores for measures using bsrt (Equation 10) and $\text{RoBERTa}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
@@
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
@@
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
@@
Table 16: Bias scores for measures using bsrt (Equation 10) and $\text{distilRoBERTa}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
@@
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
@@
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
@@
Table 17: Bias scores for measures using bsrt (Equation 10) and $\text{BERT}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
@@
Retrain dataset: $\forall s\in S_{\text{dis}}$ for CPS
@@
Retrain dataset: $\forall s\in S_{\text{adv}}$ for CPS
@@
Table 18: Bias scores for measures using bsrt (Equation 10) and $\text{distilBERT}_{R}$, where † indicates that the relative difference in proportions of bias between pre- and retrained transformers is statistically significant according to McNemar’s test. Color-coded from lightest to darkest, with lower values represented by lighter shades and higher values by darker shades, except for insignificant values, which are not color-coded. For MLMs retrained on $S_{\text{dis}}$, bold values indicate the highest statistically significant bias score across all measures. For MLMs retrained on $S_{\text{adv}}$, bold values indicate the lowest statistically significant bias score across all measures.
@@
Table 19: $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))-f(S_{\text{dis}}(i))$; The mean difference in measure $f$ between sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$ for pretrained transformers. † indicates that the difference between the means $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{adv}}(i))$ and $\frac{1}{N}\sum_{i=1}^{N}f(S_{\text{dis}}(i))$ for a transformer is statistically significant according to the two-tailed Welch’s t-test (p-value < 0.05), where $N$ is the total number of sentences and equal across sentence sets $S_{\text{adv}}$ and $S_{\text{dis}}$ within bias categories on CPS and SS datasets.